Consider a linear regression model with n predictor variables X1, . . ., Xk and a...

  1. Consider a linear regression model with n predictor variables X1, . . ., Xk and a target variable y: y= β0+β1X1+…+βkXkh6q4WdBdmvFA84EvAJBMDP2bgHyI6BMf07Cf5AAA . We take n measurements of the predictor and target variables to obtain the following matrix equation:

y=Xβ+ε5VjJtkkuQXadTmH4uxnY6Xy89vD71ffz91fVEzdWy:nx1, X:nxk+1ALHSE9Mt+CK8AAAAAElFTkSuQmCC

SSE=εTε,                   ε=y-XβsLlFrPerPZg6aCZJZcxH4BwwtVjrGqVDWAAAAAEl

Calculate the number of degrees of freedom of SSE.
Answer #1

- Bo + B. X + 2 2 X2+...+ BXE - i=1,2,...n Total degrees of freedom = df Iss Total) = n-1 No. of_predictors = kt! ob (SS Regr